Below is an ANOVA summary table for a 3x5 ANOVA reflecting the data of 12 subjects. Some of the df are incorrect, and/or it may be missing degrees of freedom.  Use the blank space below to fix the incorrect and/or provide the missing degrees of freedom (do not do more than this). What type of ANOVA design does this table represent?

Source SS Df MS F   
Subjects 11
A 2
Error A 21
B 4
Error B 11
AxB 8
Total 179
Answer #1

Here, this table represent spli-plot design ANOVA with two two factors.

ANOVA table for split-plot design is -

Source of variation SS df MS F-value
Replicate Ryy r-1 Ry Fr
A Ayy p-1 Ay Fa
Error A Eyya (r-1)(p-1) Eya Fea
B Byy q-1 By Fb
Error B Eyyb p(r-1)(q-1) Eyb -
A\timesB (AB)yy (p-1)(q-1) (AB)y Fab
Total Tyy rpq - 1 - -

Here, we have 12 subjects.

Hence, r = 12

Given ANOVA table for split-plot design is -

Source of variation SS df MS F-value
Replicate Ryy r-1 = 11 Ry Fr
A Ayy p-1 = 2 Ay Fa
Error A Eyya (r-1)(p-1) = 21 Eya Fea
B Byy q-1 = 4 By Fb
Error B Eyyb p(r-1)(q-1) = 11 Eyb -
A\timesB (AB)yy (p-1)(q-1) = 8 (AB)y Fab
Total Tyy rpq - 1 = 179 - -

From the table, we have -

r - 1 = 11 \Rightarrow r = 12

p - 1 = 2  \Rightarrow p = 3

So, (r-1)(p-1) = (11)(2) = 22

q - 1 = 4 \Rightarrow q = 5

So, p(r-1)(q-1) = (3)(11)(4) = 132

(p-1)(q-1) = (2)(4) = 8

rpq - 1 = (12)(3)(5) - 1 = 180 - 1 = 179

Hence, the corrected ANOVA table is as follows -

Source of variation SS df MS F-value
Replicate Ryy r-1 = 11 Ry Fr
A Ayy p-1 = 2 Ay Fa
Error A Eyya (r-1)(p-1) = 22 Eya Fea
B Byy q-1 = 4 By Fb
Error B Eyyb p(r-1)(q-1) = 132 Eyb -
A\timesB (AB)yy (p-1)(q-1) = 8 (AB)y Fab
Total Tyy rpq - 1 = 179 - -